OpenStudy (babyslapmafro):

Please help me evaluate the following limit by converting to spherical coordinates.

OpenStudy (babyslapmafro):

Please help me evaluate the following limit by converting to spherical coordinates: \[(\rho,\theta,\phi)\] and by observing that: \[\rho \rightarrow 0^{+ } \] if and only if: \[(x,y,z)\rightarrow(0,0,0)\]

OpenStudy (babyslapmafro):

\[\lim_{(x,y,z) \rightarrow (0,0,0)}\frac{ xyz }{ x^{2}+y ^{2}+z ^{2} }\]

OpenStudy (anonymous):

In spherical coordinates you have \[\begin{cases} x=\rho\sin\theta\cos\phi\\ y=\rho\sin\theta\sin\phi\\ z=\rho\cos\theta\\ \rho^2=x^2+y^2+z^2 \end{cases}\] So you have \[\begin{align*}\lim_{(x,y,z)\to(0,0,0)}\frac{xyz}{x^2+y^2+z^2}&=\lim_{\rho\to0^+}\frac{(\rho\sin\theta\cos\phi)(\rho\sin\theta\sin\phi)(\rho\cos\theta)}{\rho^2}\\ &=\lim_{\rho\to0^+}\rho\sin^2\theta\cos\theta \sin\phi\cos\phi \end{align*}\]